OpenStudy (anonymous):

Please need help with 1 problem! The functions f(x) and g(x) are described below: f(x) = 32x + 8 g(x) = 32x - 9 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units.

OpenStudy (anonymous):

Please, Anyone?

OpenStudy (anonymous):

Ill give metal

OpenStudy (stefrheart):

okay so when its like that what type of shifting will you be doing? up/down left/right?

OpenStudy (anonymous):

Is it going up?

OpenStudy (anonymous):

Or down?

OpenStudy (stefrheart):

now you are starting at the 8 right?

OpenStudy (asevilla5):

You might find the answer by evaluating both functions at x=0 and finding the difference by plugging in 0 for x. There is no y. What does f(0) =, plug 0 into 32x+8 Then do the same for g(0)=, 32x-9 The subtract g(0) from f(0) to get how many units you shift down. Remember what happens when you subtract a negative integer. and end up w/ f(x) =8, and g(x) = -9. it asks how many units you shift it down. At x= 0 it is at 8. For g(x) it is at -9 so you shift it down 8 to get to 0, but you still have to shift it down 9 more unites to get to -9. The total number of unites you have to shift down starting at 8 to get to -9 is 17.

OpenStudy (kl0723):

17 units down

OpenStudy (anonymous):

Thanks Guts!

OpenStudy (anonymous):

you are only concern with the vertical shift which is determined by the constant terms (y-intercepts) of the two functions. It is already mentioned from your question that g(x) is obtained by \(\Large shifting\ down\) the graph of f(x) y=mx + \(b\) , just focus on \(b\) so from the function f(x)=mx+b subtract the value of b from this function to the b of g(x)=mx+b which is 8-(-9)=17 thus, 17 units down

OpenStudy (anonymous):

Wish i could give more metals.